A man has only 5, 10 and 20 dollar bills in his purse.
If in all he has 13 bills totalling 125 dollars, find
the number of bills of each type he has. Sir Please
answer.

Answer by Edwin McCravy(20067)

He must have an odd number of 5's, otherwise
the sum would be a multiple of 10.  But he
has to have more than just 1 five, because
there is no way to make the remaining 120
with 12 bills other than with 12 10's, but
that would be using no 20's.  So

1. If he has 3 5's, he must make the remaining
110 with 10 tens and twenties.

3 fives, 9 tens, and 1 twenty. 
3+9+1 = 13
(3)(5)+(9)(10)+1(20) = 15+90+20 = 125

or

2. If he has 5 5's, he must make the remaining
100 with 8 tens and twenties.

5 fives, 6 tens, and 2 twenties.
5+6+2 = 13
(5)(5)+(6)(10)+2(20) = 25+60+40 = 125

or

3. If he has 7 5's, he must make the remaining
90 with 6 tens and twenties.

7 fives, 3 tens, and 3 twenties.
7+3+3 = 13
(7)(5)+(3)(10)+3(20) = 35+30+60 = 125

Edwin